OpenStudy (anonymous):

write an equation of a cosine function with a > 0, amplitude 3, and period pi.

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hello

OpenStudy (misty1212):

amplitude of 3 means put a 3 out front like this \[f(x)=3\cos(bx)\] now we only need \(b\)

OpenStudy (misty1212):

you want the period to be \(\pi\) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I appreciate your help. I've been stairing at this for about 40 minutes

OpenStudy (misty1212):

to find \(b\) set \[\frac{2\pi}{b}=\pi\]and solve for \(b\)

OpenStudy (misty1212):

you get \(b=2\) pretty much in your head

OpenStudy (anonymous):

yep

OpenStudy (misty1212):

final answer \[f(x)=3\cos(2x)\] that is all

OpenStudy (anonymous):

ok I was making this way harder in my head. Appreciate the help.

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]